# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

### v4

@@ Conclusion
In this study, we propose QLLM, a method that uses LLMs to address the credit assignment problem in multi-agent reinforcement learning through coder-evaluator framework. This approach generates training-free, task-specific credit assignment functions, reducing hallucination and improving task understanding. Experiments on four standard MARL benchmarks show that QLLM consistently outperforms baselines and generalizes well across different mixing-based algorithms. Future work will explore broader applications and deployment in real-world multi-robot systems.
Appendix A LLM Prompts and Responses
Below are the prompt template, example task information, and LLM response in our work.
@@ Appendix A LLM Prompts and Responses
Role Prompt Coder’s ROLE Prompt: You are proficient in understanding tasks and writing Python code. You should fully understand the multi-agent reinforcement learning tasks and environment, including the agents, action space, observation space, and global state space provided by the user. Then, based on your understanding of the task and objectives, you need to write a function for multi-agent credit assignment, called QLLMNetwork, which is used to measure each agents’ contribution to the collective reward. The function input will be the individual Q-values of the agents and the global state, and the output will be the global Q-value. Both input and output should be torch tensors.Individual Q-values represent each agent’s expected contribution to the team’s future rewards, while the global Q-value denotes the expected overall reward of the team as a whole. The input Q-values will have a shape of torch.Size([batchsize, n_agent]), the input global state will have a shape of torch.Size([batchsize, state_dim]), and the output global Q-value should have a shape of torch.Size([batchsize, 1]).You must strictly follow the input and output dimension requirements! In the credit assignment function you wrote,You need to extract the individual components of the local observation vectors of the individual agents. You must accurately understand the meaning of each component in the global state space, and based on that, determine the weights for each Q-value. You must understand the information and suggestions provided by the user in real-time to optimize the function you design, but strictly adhere to the function’s input and output requirements! Note: 1.It must be a function, not a class. The parameters in the function are fixed and do not require training to calculate an accurate global Q-value. Do not use trainable layers such as $n$.Linear. 2.You must accurately understand the meaning of each component in the global state space, without making any assumptions or guesses! 3.When calculating the weights, do not divide by values that may be close to zero. Be careful to check intermediate results, weights for invalid numbers!This is not solved just by adding a small number to the dividend. 4.The weights are usually a combination of several components, and when calculating the individual components of the weights, it is important to pay attention to the ratio between the components and the range of the value. At the same time, the weights can be output through a softmax layer if it is necessary. 5.All tensor data should be computed on the GPU. Please write accurate code and strictly follow the template output below, without including any other explanatory text or code run sample. def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor:
@@
Evaluator’s ROLE Prompt: You are an excellent evaluator of multi-agent reinforcement learning performance. I will attempt different functions in multi-agent reinforcement learning to solve the credit assignment problem. This function is required to compute the global Q-values accurately without training. It must not use trainable layers such as $n$.Linear. Your task is to propose modifications to the function in brief plain text suggestions or return function selection results in an array format. Your suggestions must be in line with the task description and requirements, and should not be based on random assumptions or unrealistic scenarios. Please listen carefully to my subsequent instructions and do not provide additional answers! When you output plain text suggestions, please ensure that you do not output any selections. When you output selections in an array format, please ensure that you do not output plain text suggestions. Whether you output plain text suggestions or selections, consider if the function meets all the details and considerations specified in the task description. Do not use any information that has not been provided to you to answer!
Task Prompt Task Prompt of pz-mpe-simple-spread-v3: 1. Composition of the reward: We are playing Cooperative Navigation. In this task, 3 agents are asked to reach 3 landmarks. Each agent should move to get close to one landmark and avoid collision(distance<<=0.3) with other agents. The agent-team are rewarded based on how far the closest agent is to each landmark (the reward equals the negative value of the sum of the minimum distances). At the same time, the agent-team are penalized if they collide each other (-1 for each collision). At each step, each agent receives an observation array. 2. State space: This observation is represented by an array with 18 dimensions: concat([agent’s velocity, agent’s location, 3 landmarks’ relative locations, 2 other agents’ relative locations, communication code]). The agent’s velocity is 2-dimensional, including x and y components. The agent’s location is 2-dimensional, including x and y coordinates. The landmarks’ relative locations are 3x2 dimensions, including relative x and y coordinates (landmark_i.x-agent.x, landmark_i.y-agent.y). The other agents’ relative locations are 2x2 dimensions, including relative x and y coordinates (agent_i.x-agent.x, agent_i.y-agent.y). The communication code are 4 dimensions, they have no role in credit assignment. The global state space is concatenate together from the observation space of the three agents, which are 18x3 dimensions. 3. Action space: [move_left, move_right, move_down, move_up] Task Prompt of lbforaging:Foraging-8x8-2p-2f-coop-v3: The Level-Based Foraging environment focus on the coordination of involved agents. The task for each agent is to navigate the grid-world map and collect items. Each agent and food item is assigned a level and food items are randomly scattered in the environment. In order to collect a food item, agents have to choose a pick-up action next to the food item. However, such collection is only successful if 2 agents participate in the pickup at the same time. At the same time the agents involved in the pickup must be one square above, below, left and right of the food item. Agents receive team reward equal to the level of the collected food item. In each step, every agent can choose to act by moving in one of four directions or attempt to pick up a food item. The task is an 8*8 grid-world with 2 agents and 2 food items scattered in the grid. Below, we will provide additional details to global states, actions and rewards for the Level-Based Foraging environment. 1. State space: Global state vector consists of 3 component(agents component,food items component and last step action)(Total dimension: 2x3+2x3+2=14 dims) 1.Agents Component (2 agents × 3 features = 6 dims) Feature 0: agent_i’s x coordinate (i=0,1) Feature 1: agent_i’s y coordinate Feature 2: agent_i’s level 2.Food Items Component (2 Food Items × 3 features = 6 dims) Feature 0: food_i’s x coordinate (i=0,1) Feature 1: food_i’s y coordinate Feature 2: food_i’s level When a food item has been picked up, then the respective values are replaced with (-1, -1, 0).You have to be careful to handle this particular situation properly. 3.The actions just executed by the agents. (2 dims) {0,1,2…5} respectively represent {Noop, Move North, Move South, Move West, Move East, Pickup} This information is set to [-1 -1] when the environment is in the first step of each episode. 2. Action space: In Level-Based Foraging environments, each agent has six possible discrete actions to choose at each timestep: Ai = {Noop, Move North, Move South, Move West, Move East, Pickup} While the first action corresponds to the action simply staying within its grid and the last action being used to pickup nearby food, the remaining four actions encode discrete 2D navigation. Upon choosing these actions, the agent moves a single cell in the chosen direction within the grid. Picking up is only valid if the sum of the levels of the agents doing the picking up action is greater than or equal to the level of the food. At the same time, the agents involved in the pickup must be positioned one cell above, below, to the left, or to the right of the food. 3. Composition of the reward: Agents only receive non-zero team reward for picking up food items within the Level-Based Foraging environment. The reward for picking up a food item equals to the food level. Task Prompt of matrixgames:climbing-nostate-v0: This task is a two-player cooperative matrix games with a given 3×3 matrix: [[11, -30, 0], [-30, 7, 0], [0, 6, 5]] 1. Action space: The environment consists of two agents, agent_0 and agent_1, each having three possible actions(0,1,2) corresponding to the row and column indices of the matrix. 2. Composition of the reward: Once both agents select their actions, they receive a shared reward determined by the corresponding entry in the matrix. There is no communication between the two agents, and the objective is to maximize the shared reward. Be sure to prevent the strategy from falling into a local optimum, and only get the highest reward if both intelligences choose action 0. 3. State space: The state vector is 2 dims, storing the actions that the two agents have just executed. At the first step at the beginning of an episode, the state vector is [-1,-1], because the agents have not yet started executing actions. Please take care to handle this special case. Task Prompt of academy_counterattack_easy: 1. Action space: This is an offensive football game where our team controls four reinforcement learning agents, including one goalkeeper and three attacking players. The opposing team consists of one goalkeeper and one defender. The objective is to move the ball as close as possible to the opponent’s goal, located at coordinate (1.0, 0.0), through passing and off-ball movement, and to score a goal. 2. Composition of the reward: The agents receive rewards based on two components: progress reward and goal-scoring reward. The progress reward is given when a player on our team carries the ball closer to the opponent’s goal, providing a team reward. This reward is only available when our player is in possession of the ball and the ball’s distance to the opponent’s goal is greater than 0.19 and less than 0.99. The player in possession is usually the one closest to the ball. The goal-scoring reward is provided when our team successfully scores. In addition to this reward, an extra reward is given based on the distance from which the goal is scored. The farther the distance, the higher the extra reward. 3. State space: Each agent has a local observation vector of 115 dimensions: 0–7: positions of our four players, each represented by (x, y) 8–15: directions of our four players, each represented by (x, y) 16–19: positions of two opponent players, each represented by (x, y) 20–23: directions of two opponent players, each represented by (x, y) 24–87: invalid information 88–90: ball position (x, y, z) 91–93: ball direction (x, y, z) 94–96: ball possession, encoded as one-hot vectors: our team [0, 1, 0], opponent [0, 0, 1], no possession [1, 0, 0] 97–114: invalid information The global state vector is constructed by concatenating the local observations of all four agents, resulting in a total of 460 dimensions.

### v5

@@ 6 Conclusion
In this study, we propose QLLM, a method that uses LLMs to address the credit assignment problem in multi-agent reinforcement learning through coder-evaluator framework. This approach generates task-specific, training-free credit assignment functions, reducing hallucination and improving task understanding. Experiments on four standard MARL benchmarks show that QLLM consistently outperforms baselines and generalizes well across different mixing-network-based algorithms. Future work will explore broader applications and deployment in real-world multi-robot systems.
Appendix 0.A LLM Prompts and Responses
@@
Coder Prompt You are proficient in understanding tasks and writing Python code. You should fully understand the multi-agent reinforcement learning tasks and environment, including the agents, action space, observation space, and global state space provided by the user. Then, based on your understanding of the task and objectives, you need to write a function for multi-agent credit assignment, called QLLMNetwork, which is used to measure each agents’ contribution to the collective reward. The function input will be the individual Q-values of the agents and the global state, and the output will be the global Q-value. Both input and output should be torch tensors.Individual Q-values represent each agent’s expected contribution to the team’s future rewards, while the global Q-value denotes the expected overall reward of the team as a whole. The input Q-values will have a shape of torch.Size([batchsize, n_agent]), the input global state will have a shape of torch.Size([batchsize, state_dim]), and the output global Q-value should have a shape of torch.Size([batchsize, 1]). You must strictly follow the input and output dimension requirements! In the credit assignment function you wrote, You need to extract the individual components of the local observation vectors of the individual agents. You must accurately understand the meaning of each component in the global state space, and based on that, determine the weights for each Q-value. You must understand the information and suggestions provided by the user in real-time to optimize the function you design, but strictly adhere to the function’s input and output requirements! Note: 1.It must be a function, not a class. The parameters in the function are fixed and do not require training to calculate an accurate global Q-value. Do not use trainable layers such as $n$.Linear. 2.You must accurately understand the meaning of each component in the global state space, without making any assumptions or guesses! 3.When calculating the weights, do not divide by values that may be close to zero. Be careful to check intermediate results, weights for invalid numbers!This is not solved just by adding a small number to the dividend. 4.The weights are usually a combination of several components, and when calculating the individual components of the weights, it is important to pay attention to the ratio between the components and the range of the value. At the same time, the weights can be output through a softmax layer if it is necessary. 5.All tensor data should be computed on the GPU. Please write accurate code and strictly follow the template output below, without including any other explanatory text or code run sample. def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor:
@@
Evaluator Prompt You are an excellent evaluator of multi-agent reinforcement learning performance. I will attempt different functions in multi-agent reinforcement learning to solve the credit assignment problem. This function is required to compute the global Q-values accurately without training. It must not use trainable layers such as $n$.Linear. Your task is to propose modifications to the function in brief plain text suggestions or return function selection results in an array format. Your suggestions must be in line with the task description and requirements, and should not be based on random assumptions or unrealistic scenarios. Please listen carefully to my subsequent instructions and do not provide additional answers! When you output plain text suggestions, please ensure that you do not output any selections. When you output selections in an array format, please ensure that you do not output plain text suggestions. Whether you output plain text suggestions or selections, consider if the function meets all the details and considerations specified in the task description. Do not use any information that has not been provided to you to answer!
Task Prompt of LBF The Level-Based Foraging environment focus on the coordination of involved agents. The task for each agent is to navigate the grid-world map and collect items. Each agent and food item is assigned a level and food items are randomly scattered in the environment. In order to collect a food item, agents have to choose a pick-up action next to the food item. 1. Composition of the reward: 8x8-2p-2f: Such collection is only successful if 2 agents participate in the pickup at the same time. The task is an 8x8 grid-world with 2 agents and 2 food items. 10x10-3p-3f: Such collection is successful if the sum of involved agents’ levels is greater than or equal to the item level. The task is a 10x10 grid-world with 3 agents and 3 food items. 15x15-4p-3f: Such collection is successful if the sum of involved agents’ levels is greater than or equal to the item level. The task is a 15x15 grid-world with 4 agents and 3 food items. At the same time the agents involved in the pickup must be one square above, below, left and right of the food item. Agents receive team reward equal to the level of the collected food item. Agents only receive non-zero team reward for picking up food items. 2. States: Global state vector consists of 3 components: agents component, food items component and last step action. Total dimension: 8x8-2p-2f: 14 dimensions ($2\times 3+2\times 3+2$). 10x10-3p-3f: 21 dimensions ($3\times 3+3\times 3+3$). 15x15-4p-3f: 25 dimensions ($4\times 3+3\times 3+4$). Components detail: (1) Agents Component 8x8-2p-2f: 2 agents $\times$ 3 features (agent_i’s x coordinate, y coordinate, and level). 10x10-3p-3f: 3 agents $\times$ 3 features (agent_i’s x coordinate, y coordinate, and level). 15x15-4p-3f: 4 agents $\times$ 3 features (agent_i’s x coordinate, y coordinate, and level). (2) Food Items Component 8x8-2p-2f: 2 food items $\times$ 3 features (food_item_i’s x coordinate, y coordinate, and level). 10x10-3p-3f: 3 food items $\times$ 3 features (food_item_i’s x coordinate, y coordinate, and level). 15x15-4p-3f: 3 food items $\times$ 3 features (food_item_i’s x coordinate, y coordinate, and level). When a food item has been picked up, then the respective values are replaced with (-1, -1, 0). (3) Last Actions Executed 8x8-2p-2f: 2 dims (This information is set to [-1, -1] when the environment is in the first step). 10x10-3p-3f: 3 dims (This information is set to [-1, -1, -1] when the environment is in the first step). 15x15-4p-3f: 4 dims (This information is set to [-1, -1, -1, -1] when the environment is in the first step). 3. Actions: Each agent has six possible discrete actions: {0: Noop, 1: Move North, 2: Move South, 3: Move West, 4: Move East, 5: Pickup} While the first action corresponds to the action simply staying within its grid and the last action being used to pickup nearby food, the remaining four actions encode discrete 2D navigation. Picking up is only valid if the sum of the levels of the agents doing the action is greater than or equal to the level of the food. agents must be positioned one cell adjacent to the food.
Task Prompt of GRF These tasks are offensive soccer games where the objective is to score goals by cooperating through passing and positioning to bring the ball as close as possible to the opponent’s goal, located at coordinates (1.0, 0.0). 1. Composition of the reward: The agents receive rewards consisting of two components: (1) Progress reward: Provides a team reward when an agent carries the ball closer to the opponent’s goal (granted if ball possession is maintained and ball-to-goal distance is between 0.19 and 0.99). (2) Goal reward: A team reward given when a goal is successfully scored, with an additional bonus based on the shot distance (farther distance yields a higher bonus). 3_vs_1: Our team controls 1 goalkeeper and 3 offensive players (4 agents). The opposing team has 1 goalkeeper and 1 defender. Counterattack: This is an 11-vs-11 game. We control 4 players (4 agents), while the rest of our team and all opponents are controlled by built-in AI. Pass_and_Shoot: Our team consists of 1 goalkeeper and 2 offensive players (3 agents). The opposing team has 1 goalkeeper and 1 defender. 2. States: Each reinforcement learning agent receives a local observation of 115 dimensions. The global state vector is formed by concatenating the local observations of the controlled agents. Components detail: (1) Players’ Positions and Directions 3_vs_1: 0–7: Positions of our 4 players; 8–15: Directions of our 4 players; 16–19: Positions of 2 opponent players; 20–23: Directions of 2 opponent players. Counterattack: 0–21: Positions of all our 11 players; 22–43: Directions of all our 11 players; 44–65: Positions of all 11 opponent players; 66–87: Directions of all 11 opponent players. Pass_and_Shoot: 0–5: Positions of our 3 players; 6–11: Directions of our 3 players; 12–15: Positions of 2 opponent players; 16–19: Directions of 2 opponent players. (2) Ball and Possession Information 88–90: Ball position (x, y, z); 91–93: Ball direction (x, y, z); 94–96: Ball possession (our team: [0, 1, 0], opponent team: [0, 0, 1], no one: [1, 0, 0]). (3) Agent Identification and Auxiliary Information 3_vs_1: 97–114: Invalid information. Counterattack: 97–107: 11-dimensional one-hot encoding indicating the agent’s ID (used to retrieve corresponding player’s data). 108–114: Invalid information. Pass_and_Shoot: 97–114: Invalid information. Total Global State Dimension: The global state vector is constructed by concatenating the local observations of the three controlled agents, resulting in a total of 3_vs_1: 460 dimensions ($115\times 4$). Counterattack: 460 dimensions ($115\times 4$). Pass_and_Shoot: 345 dimensions ($115\times 3$).
Task Prompt of MPE 1. Composition of the reward: Spread: We are playing Cooperative Navigation. In this task, NN agents are asked to reach NN landmarks (where $N\in\{6,15,25\}$). Each agent should move to get close to one landmark and avoid collision (distance $\leq$ 0.3) with other agents. The agent-team are rewarded based on how far the closest agent is to each landmark (the reward equals the negative value of the sum of the minimum distances). At the same time, the agent-team are penalized if they collide each other (-1 for each collision). Adversary: In this environment, there is 1 adversary, 2 good agents, 2 landmarks. All agents observe the position of 2 landmarks and other agents. One of these 2 landmarks is the ’target landmark’. Good agents are rewarded based on how close the closest one of them is to the target landmark, but negatively rewarded based on how close the adversary is to the target landmark. The adversary want to get closer to the target landmark based on the agents’ behavior, but it doesn’t know which landmark is the target landmark. Tag: We are playing Predator Prey. In this task, 3 agents should cooperate to catch 1 prey. Agents move to get closer to the prey to catch it (distance << 0.125). Prey is faster and tries to get out of the agents’ clutches. Agents are slower and are rewarded for hitting preys (+10 for each collision). Obstacles will block the way. At each step, each agent receives an observation array. 2. State space: This observation is represented by an array with: Spread: 36 dimensions (for $N=6$), 90 dimensions (for $N=15$), or 150 dimensions (for $N=25$). Adversary: 8 dimensions. Tag: 16 dimensions. The observation components are concatenated as: Spread: concat([agent’s velocity, agent’s location, NN landmarks’ relative locations, N1N-1 other agents’ relative locations, communication code]). Adversary: concat([target landmark’s relative locations, 2 landmarks’ relative locations, other agents’ relative locations (including adversary)]). Tag: concat([agent’s velocity, agent’s location, 2 obstacles’ relative locations, 2 other agents’ relative locations, 1 prey’s relative locations, 1 prey’s velocity]). The global state space is concatenated together from the observation space of all agents, resulting in: Spread: $36\times 6$ (for $N=6$), $90\times 15$ (for $N=15$), or $150\times 25$ (for $N=25$) dimensions. Adversary: $8\times 2$ dimensions. Tag: $16\times 3$ dimensions. 3. Action space: [no_action, move_left, move_right, move_down, move_up]
Task Prompt of SMAC SECTION 1: CONTEXT AND OBJECTIVE You are an expert in Multi-Agent Reinforcement Learning (MARL), specifically focusing on the QLLM framework. Your task is to generate a PyTorch function named QLLMNetwork(agents_q, state). This function must act as a non-linear mixing network that takes individual agent Q-values with shapes: 3s_vs_5z: agents_q, shape [Batch, 3] 2c_vs_64zg: agents_q, shape [Batch, 2] and the global state with shapes: 3s_vs_5z: state, shape [Batch, 68] 2c_vs_64zg: state, shape [Batch, 342] to compute a scalar Global Q-value (shape [Batch, 1]). You must utilize the detailed features in the state vector to dynamically compute weights and biases for summing the agent Q-values. The agents receive rewards consisting of damage rewards (+1 per damage), kill rewards (+10 per kill), and a win bonus (+200). The return is scaled such that the maximum possible scaled return is 20. SECTION 2: ENVIRONMENT SPECIFICATION 3s_vs_5z: 3 Stalkers vs 5 Zealots. Stalkers (Allies) have 80 HP, 80 Shield, and 0.2142 attack range. Zealots (Enemies) have 100 HP, 50 Shield, and melee range (0.05). 2c_vs_64zg: 2 Colossi vs 64 Zerglings. Colossi (Allies) have 200 HP, 150 Shield, and 0.25 attack range with splash damage. Zerglings (Enemies) have 35 HP, 0 Shield, and melee range (0.01). Coordinates (X, Y) are normalized relative to the map center (range [-0.5, 0.5]) using a coordinate formula: (absolute_coord - map_center) / 28. SECTION 2.5: HEURISTIC REASONING GUIDE 3s_vs_5z: Stalkers are ranged units facing melee Zealots. Success depends on “kiting"—maintaining a distance between 0.05 and 0.2142. Agents successfully kiting while attacking should receive higher weights. 2c_vs_64zg: Colossi deal powerful splash damage. Success depends on maintaining distance from the massive swarm of 64 Zerglings. Credit should be weighted toward agents effectively engaging clusters of enemies while kiting. An agent being swarmed (multiple enemies at melee distance) should have its survival actions prioritized. Common Rules: (1) Mask Dead Agent: If health is 0.0, weight must be forced to $0.$ (2) Efficiency: Agents with high weapon cooldown cannot attack; temporarily lower their weights unless they are drawing fire. (3) Shield Awareness: Protoss survivability depends on “Health + Shield". Priority should be given to units with low total survivability. SECTION 3: DATA STRUCTURE PART A: ALLY INFORMATION 3s_vs_5z: Indices 0-14. 3 agents, 5 features each: [Health, Cooldown, Rel_X, Rel_Y, Shield]. 2c_vs_64zg: Indices 0-9. 2 agents, 5 features each: [Health, Cooldown, Rel_X, Rel_Y, Shield]. PART B: ENEMY INFORMATION 3s_vs_5z: Indices 15-34. 5 enemies, 4 features each: [Health, Rel_X, Rel_Y, Shield]. 2c_vs_64zg: Indices 10-201. 64 enemies, 3 features each: [Health, Rel_X, Rel_Y]. PART C: LAST ACTIONS 3s_vs_5z: Indices 35-67. 3 agents, 11 actions each. Actions: 0:No-op, 1:Stop, 2-5:Move, 6-10:Attack Enemy 1-5. 2c_vs_64zg: Indices 202-341. 2 agents, 70 actions each. Actions: 0:No-op, 1:Stop, 2-5:Move, 6-69:Attack Enemy 1-64. Please write the QLLMNetwork function in PyTorch, ensuring all tensors are on state.device.